# Further author information: (Send correspondence to Songjiang-Lai.)

### v1

@@ 2 HYBRID ATTENTION AGGREGATION TRANSFORMER
Figure 1 illustrates the comprehensive structure of HAAT. SDRCB incorporates Swin Transformer Layers and transition layers into each Residual Deep feature extraction Group (RDG), enhancing the receptive field while using fewer parameters and a more streamlined design, resulting in superior performance. Furthermore, we provide HGAB to describe cross-area similarity for enhanced picture reconstruction. The architecture of HGAB, seen in Fig. 2, comprises a Mix Attention Layer (MAL) and a Multi-Layer Perceptron (MLP) layer. HGAB employs sparse self-attention to augment global feature interactions while controlling computational complexity, facilitating the joint modeling of analogous features for enhanced picture reconstruction. Furthermore, the employed channel attention mechanism can help the model extract more effective information between different channels.
@@ 2.2 Hybrid Grid Attention Block(HGAB)
The GAB consists of a Mix Attention Layer (MAL) and an MLP layer. Regarding the MAL, we first split the input feature $F_{in}$ into two parts by channel: $F_{G}\in R^{(H\times W\times{C/2})}$ and $F_{W}\in R^{(H\times W\times{C/2})}$. Besides, feed the input to another branch to perform channel attention operation. Subsequently, we split $F_{W}$ into two parts by channel again and input them into W-MSA and SW-MSA, respectively. Meanwhile, $F_{G}$ is input into Grid-MSA[9].The computation process of MAL is as follows:
@@
where $X_{W_{1}}$, $X_{W_{2}}$, $X_{G}$ and $X_{C}$ are the output features of W-MSA, SW-MSA, Grid-MSA, and CA, respectively. Furtheremore, it should be noted that we adopt the post-norm method in GAB to enhance the network training stability. For a given input feature $F_{in}$, the computation process of HGAB is:
@@ 4 CONCLUSION
Acknowledgements.
This unnumbered section is used to identify those who have aided the authors in understanding or accomplishing the work presented and to acknowledge sources of funding [Lamport94, Alred03, Goossens97].

### v2

@@ 2 HYBRID ATTENTION AGGREGATION TRANSFORMER
Figure 1 illustrates the comprehensive structure of HAAT. SDRCB incorporates Swin Transformer Layers and transition layers into each Residual Deep feature extraction Group (RDG), enhancing the receptive field while using fewer parameters and a more streamlined design, resulting in superior performance. Furthermore, we provide HGAB to describe cross-area similarity for enhanced picture reconstruction. The architecture of HGAB, seen in Figure 2, comprises a Mix Attention Layer (MAL) and a Multi-Layer Perceptron (MLP) layer. HGAB employs sparse self-attention to augment global feature interactions while controlling computational complexity, facilitating the joint modeling of analogous features for enhanced picture reconstruction. Furthermore, the employed channel attention mechanism can help the model extract more effective information between different channels.
@@ 2.2 Hybrid Grid Attention Block(HGAB)
The structure of Hybrid Grid Attention Block (HGAB) is illustrated in Figure 2. The HGAB consists of a Mix Attention Layer (MAL) and an MLP layer. Regarding the MAL, we first split the input feature $F_{in}$ into two parts by channel: $F_{G}\in R^{(H\times W\times{C/2})}$ and $F_{W}\in R^{(H\times W\times{C/2})}$. Besides, feed the input to another branch to perform channel attention operation. Subsequently, we split $F_{W}$ into two parts by channel again and input them into W-MSA and SW-MSA, respectively. Meanwhile, $F_{G}$ is input into Grid-MSA[9].The computation process of MAL is as follows:
@@
where $X_{W_{1}}$, $X_{W_{2}}$, $X_{G}$ and $X_{C}$ are the output features of W-MSA, SW-MSA, Grid-MSA, and CA, respectively. Furtheremore, it should be noted that we adopt the post-norm method in HGAB to enhance the network training stability. For a given input feature $F_{in}$, the computation process of HGAB is: